OpenStudy (anonymous):

okay first i need to know what the formula is and how it works.(warning im tried so i may act silly or it may take me time to understand it:P) Solve the following using the quadratic formula. Round to the nearest tenth, if necessary. x2 + 5x - 2 = 0 {-2, 5} {-5.4, 0.4} {-1, 0.5} No real solution

OpenStudy (anonymous):

@timo86m

jimthompson5910 (jim_thompson5910):

do you know the quadratic formula?

OpenStudy (anonymous):

no sir

jimthompson5910 (jim_thompson5910):

the quadratic formula is \[\large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

in this case, a = 1, b = 5 and c = -2 plug those 3 values in and simplify

OpenStudy (anonymous):

\[x=-5+\sqrt{5^2-4*1*(-2)}/ 2*1\]

jimthompson5910 (jim_thompson5910):

close, don't forget about the plus/minus \[\Large x = \frac{-5\pm\sqrt{5^2-4(1)(-2)}}{2(1)}\]

jimthompson5910 (jim_thompson5910):

keep going and simplify

OpenStudy (anonymous):

would i go by the PEMDAS to simplfy this??

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

okay do i simplfy inside the root first?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

(-5.4,0.4)

jimthompson5910 (jim_thompson5910):

very good, you got the two solutions

OpenStudy (anonymous):

thank you for your help and the formula

jimthompson5910 (jim_thompson5910):

you're welcome